OpenStudy (anonymous):

Score: A quiz has 1 true-false question and 2 multiple choice questions that each have 4 answers. What is the probability that someone guessing all 3 answers will get 100% on the quiz?

OpenStudy (anonymous):

the probability of getting a T-F answer correct by guessing is 1/2 The probability of getting a M.C. answer correct by guessing is 1/4 When combining probabilities they are multiplied so, P = (1/2)(1/4)(1/4) = 1/32

OpenStudy (anonymous):

thanks man,